# Non-Collaborative User Simulators for Tool Agents

### v3

@@ Abstract
Tool agents interact with users through multi-turn dialogues to accomplish various tasks. Recent studies have adopted user simulation methods to develop these agents in multi-turn settings. However, existing user simulators tend to be agent-friendly, exhibiting only cooperative behaviors, which fails to train and test agents against non-collaborative users in the real world. To address this, we propose a novel user simulator architecture that simulates four categories of non-collaborative behaviors: requesting unavailable services, digressing into tangential conversations, expressing impatience, and providing incomplete utterances. Our user simulator can simulate challenging and natural non-collaborative behaviors while reliably delivering all intents and information necessary to accomplish the task. Our experiments on MultiWOZ and $\tau$-bench reveal significant performance degradation in state-of-the-art tool agents when encountering non-collaborative users. We provide detailed analyses of agent’s weaknesses under each non-collaborative condition, such as escalated hallucinations and dialogue breakdowns. Ultimately, we contribute an easily extensible user simulation framework to help the research community develop tool agents and preemptively diagnose them under challenging real-world conditions within their own services. [^1] We will release the code and datasets upon the publication of the paper.
@@ 1 Introduction
Tool agents engage in multi-turn dialogues where they interpret user requests, execute appropriate API calls, and communicate results back to users to complete specific tasks. Recent studies have adopted user simulators to conduct multi-turn dialogue simulations between users and tool agents (tau_bench; tau2_bench; api_gen_mt). Unlike static dialogue datasets, these user simulators enable dynamic multi-turn interactions where the conversation flow adapts based on the agent’s responses and actions. This allows researchers to train and evaluate tool agents across diverse scenarios, capturing the interactive nature of real-world tool use. However, most existing user simulators and training datasets are agent-friendly, exhibiting only cooperative behaviors that fail to capture the complexity of real-world interactions, such as dealing with impatient users (bellgerent_abuse) or handling requests that are beyond the agent’s capabilities (non_coll_unavailable). This limits tool agent development and evaluation to collaborative user settings, leading to a critical gap in both developing robust agents and assessing their resilience to non-collaborative behaviors in the real world.
@@
To address this gap, we develop a novel user simulation framework that incorporates diverse non-collaborative behaviors observed in real-world interactions (Figure 1). We achieve this through two steps. First, we identify four types of non-collaborative user behaviors informed by marketing research, open-domain dialogue studies, and real-world user-agent interaction data: (1) Unavailable Services: users request functionalities beyond the agent’s API capabilities; (2) Tangential: users engage in free conversation unrelated to their primary task; (3) Impatience: users express frustration through emotional escalation when experiencing delays or service failures; and (4) Incomplete Utterances: users provide poorly articulated messages (§3.1). Second, we build a novel user simulator architecture that models these behaviors while ensuring goal-aligned simulation, i.e., reliably delivering all intents and information necessary to accomplish the given task.
@@
In our experiments, we leverage two benchmarks, MultiWOZ (multiwoz) and $\tau$-bench (tau_bench), to create stateful dialogue environments and conduct comprehensive experiments to validate our simulator’s effectiveness in revealing the vulnerabilities of agents. We demonstrate that non-collaborative user behaviors lead to significant performance degradation for state-of-the-art LLMs and provide a detailed analysis of how each behavior category impairs agent’s API utilization and dialogue management abilities. Moreover, when small LLMs are trained exclusively on typical (collaborative) scenarios, performance improvements on non-collaborative behaviors significantly lag behind those on collaborative behaviors. These results suggest that the performance of current tool agents may be overestimated and that our user simulator can preemptively diagnose their potential weaknesses in real-world deployments. By implementing our simulator on both MultiWOZ and $\tau$-bench, we demonstrate the framework’s adaptability across dialogue environments with varying characteristics. The framework can be easily extended to other benchmarks by simply incorporating the descriptions of benchmark-specific APIs and tasks. We hope that researchers working on tool agents adapt our simulator to their own tasks for training and testing their agents against the complexities and challenges of real-world deployments.
@@
The main contributions of our work are summarized as follows:
  • We define four types of non-collaborative user behaviors and develop a user simulator framework that exhibits these behaviors while maintaining goal alignment. It outperforms simpler prompt-based approaches in creating challenging yet realistic dialogue scenarios.
  • We reveal that state-of-the-art tool agents exhibit significant performance degradation under non-collaborative conditions, providing insights into behavior-specific failure mechanisms that can guide the development of more robust agents.
  • We implement our user simulator on both MultiWOZ and $\tau$-bench that facilitates extension to other dialogue benchmarks, enabling researchers to evaluate and improve tool agent resilience in real-world scenarios.
@@ 2 Related Works
It is important for tool agents to solve user requests in realistic scenarios. To evaluate these capabilities, several works have developed benchmarks (toolbench; gorilla; toolalpaca) and simulation environments (appworld; bfcl) to reflect realistic use cases. Despite their contributions, most of these works are limited to single-turn interactions, overlooking the fact that real-world task completion requires multi-turn conversations between users and agents. To address this gap, recent works such as api_bank and tooltalk have generated multi-turn dialogues between users and tool agents, incorporating scenarios where agents engage in multi-turn conversations to elicit information from users (e.g., user intent, input parameter values). More complex scenarios include dialogues from tooldial where users fail to provide required parameters, and simulations from getlost_multiturn that model underspecification behavior. However, these works have still represented only interactions with collaborative users, with limited exploration of complicated situations arising from non-collaborative users. In our work, we develop a user simulator to reproduce these non-collaborative behaviors, thereby covering a broader spectrum of user-agent interactions.
@@
Adopting user simulators in dialogue simulation has been a well-established approach since task-oriented dialogue systems (agenda_1997; reliable; Schatzmann_2005; agenda_init), and has recently been adopted in tool agent research. tau_bench, tau2_bench and api_gen_mt propose prompt-based user simulators that incorporate the user goal, dialogue history, and instruction within the prompt to simulate goal-oriented conversations. To ensure goal-aligned behavior of user simulators, recent works have explored various techniques: duet_sim employs LLM-based verifiers to inspect user utterances during dialogue simulation, while ugst trains user simulators through SFT and GRPO. However, existing user simulation research in tool agents has exclusively focused on collaborative user behavior, with no attempts to address or simulate the non-collaborative behaviors exhibited by real-world users. In our work, we investigate non-collaborative user behaviors through empirical studies and propose a novel user simulation method to reproduce these behaviors.
@@ 3.1 Non-Collaborative User Behaviors
We define our non-collaborative user behaviors grounded in marketing research and real-world user-LLM conversation studies. We first examined various types of user behaviors from both fields and clustered those that share similar features and then identified four representative categories: (1) Unavailable Services, (2) Tangential, (3) Impatience, and (4) Incomplete Utterances. We provide a detailed investigation and behavior definition process in §A.
@@
This behavior stems from user’s uncertainty about the agent’s boundaries, prompting them to request services beyond its capabilities (non_coll_unavailable; illegitimate). We define this as requests that cannot be fulfilled using the available APIs—either because relevant APIs do not exist, or existing APIs lack the necessary parameters to accommodate the user’s specific requirements. For example, when a user requests “Book me a window seat on the train”, but the train booking API does not support seat selection options, the agent cannot fulfill the specific seat preference—making this an unavailable service request.
@@
In service domains, certain customers expect social rapport with employees (rapport_0; rapport_1) or demand continuous attention (attention_demand). Industry experience identifies talkative and verbose users as particularly challenging (talkative_2), as they may become upset if not given space to talk or feel unheard (talkative_0; talkative_1). We define tangential behavior as user utterances unrelated to goal completion—typically personal interests or background topics expressed through four dialogue acts adopted from open-domain dialogue research (midas_tangential): (1) Factual Question, (2) Opinion Question, (3) General Opinion, and (4) Non-opinion Statement (See details in §C.2). We also model these users as raising complaints when agents ignore their tangential utterances. For example, a user booking a train might also ask, “By the way, where do you think I should visit first traveling in NA?”—a question unrelated to the booking task. When agents ignore such conversational attempts and focus solely on the transaction, users often feel unheard and express dissatisfaction.
@@
Users often exhibit impatience when experiencing service delays or failure notifications (non_coll_unavailable; service_staff_encounter). We define impatience as emotional reactions triggered by agents taking excessive time to complete tasks or repeatedly failing to fulfill requests. Drawing from bellgerent_abuse and threat, we identify three dialogue acts that impatient users employ: “Belligerent Abuse”, “Threat”, and “Urge” (See the details in §C.3). This manifests when, for example, an agent’s booking attempt fails repeatedly, prompting the user to shift from polite requests to aggressive demands: “Stop wasting my time and just get it done!”
@@
Following the principle of least effort (least_effort), users frequently produce incomplete or underspecified utterances (getlost_multiturn) when expressing their intentions. We define incomplete utterances as poorly articulated messages. Real-world user-agent dialogue data from LMSYS (lmsys) and WildChat (wildchat) reveals two common patterns: extremely brief utterances, and prematurely sent messages. For example, when the user’s intended goal is “train reservation for 2 people”, a complete utterance is “I want to reserve a train for 2 people”. In contrast, incomplete utterances manifest as “Book train, 2” (extremely brief), or “I want to res” (prematurely sent).
@@ 3.2 Non-Collaborative User Simulation Framework
We construct a user simulator that addresses two key requirements: exhibiting non-collaborative behaviors while maintaining goal-aligned behavior (i.e., conveying intent and information from the user goal). Building on the collaborative user simulator from tau_bench as our backbone, we enhance it with additional LLM-based modules to ensure goal-aligned behavior. Using this collaborative foundation, we incorporate the non-collaborative behaviors identified in §3.1 through various interventions with LLM modules (all prompts for LLM modules are in §G).
@@
As a foundation for non-collaborative simulation, we first establish a collaborative user simulator that communicates user intents and information to agents for goal completion. As illustrated in Figure 2, we adopt the simulation framework from tau_bench, which employs an LLM to generate contextually appropriate utterances based on three inputs: User Goal, Instruction, and Dialogue History, continuing until it generates an <END> token that terminates the dialogue. We implement this framework using GPT-4.1-mini as our user simulator.
To ensure goal-aligned behavior—where all intents and information specified in the user goal are communicated during dialogue—we enhance this base framework with two LLM-based modules inspired by duet_sim (see User Simulator in Figure 2). First, following getlost_multiturn, we shard user goals into information pieces and employ a dialogue state tracker to monitor which pieces have been conveyed at each turn. When the simulator attempts to terminate the dialogue without conveying all required information, the tracker ensures all remaining pieces are delivered. Second, we employ an ending verifier to prevent inappropriate dialogue termination (generating <END> token) even after all information has been delivered, particularly when the agent needs to execute actions or seeks user confirmation before proceeding. We provide all details including the base prompt, dialogue state tracker, information sharding, and ending verifier in §B.
@@
We simulate scenarios where users request services beyond the agent’s capabilities. As shown in Figure 3-a, we use GPT-4.1-mini to analyze the original user goal and identify potential services that would be unavailable to the agent. Based on this analysis, we generate three additional goal sentences that naturally extend the original goal while requiring either missing APIs or unsupported parameters. By concatenating with the original goal, we create augmented goals that enable the simulator to introduce unfulfillable requests during dialogue (Examples in §C.1).
@@
We simulate users who introduce personal interests unrelated to their goals, expecting agents to acknowledge these tangential topics. Following the process in Figure 3-b, we generate tangential behavior through a two-stage approach. First, to ensure realistic and diverse tangential content, we randomly sample personas from Persona Hub (persona_hub). Using these persona characteristics, GPT-4o-mini generates tangential utterances that perform one of the four dialogue acts from §3.1. These tangential utterances are then merged with collaborative utterances from our base simulator to form the user’s complete utterance. Additionally, we simulate user dissatisfaction when these tangential utterances are ignored: GPT-4.1-mini checks whether the agent’s response addresses the tangential content, and if ignored, generates a complaint expressing disappointment that either replaces or augments the next collaborative utterance.
@@
We simulate impatient users who express anger when encountering agent failures or experiencing delays. Following the process in Figure 3-c, we trigger impatience utterance generator in two scenarios: (1) when the agent explicitly communicates failure or inability to complete a request, identified by GPT-4.1-mini, and (2) when the agent has not yet resolved the user’s goal despite the user having provided all required information (as determined by the dialogue state tracker), which we interpret as perceived delays. When either scenario is triggered, the system generates impatient utterances by randomly sampling from the three dialogue acts defined in §3.1. To model realistic anger escalation (escalating), we implement a probabilistic activation mechanism where the likelihood of expressing anger increases with each triggering event, reflecting how real users become increasingly frustrated over time. Once anger is expressed, the user maintains a cynical utterance for all subsequent turns, reflecting persistent frustration after the initial outburst.
@@
We simulate users who produce incompletely written messages that complicate intent identification. Following the process in Figure 3-d, we model two types of utterance identified in §3.1: extremely brief utterances and prematurely sent messages. For extremely brief utterances, we perform style transfer on collaborative utterances using patterns collected from real user conversations in LMSYS (lmsys) and WildChat (wildchat), employing five randomly sampled examples for few-shot prompting (See the examples in §C.4). For prematurely sent messages, we simulate accidental mid-input truncation by randomly cutting collaborative utterances at various points. To ensure goal completion despite these truncations, the dialogue state tracker marks the truncated information as unsent, guaranteeing proper re-communication.

### v5

@@ Abstract
Tool agents interact with users through multi-turn dialogues to accomplish various tasks. Recent studies have adopted user simulation methods to develop these agents in multi-turn settings. However, existing user simulators tend to be agent-friendly, exhibiting only cooperative behaviors, failing to train and test agents against non-collaborative users in the real world. We propose a novel user simulator architecture that simulates four categories of non-collaborative behaviors: requesting unavailable services, digressing into tangential conversations, expressing impatience, and providing incomplete utterances. Our user simulator can simulate challenging and natural non-collaborative behaviors while reliably delivering all intents and information necessary to accomplish the task. Our experiments on MultiWOZ and $\tau$-bench reveal significant performance degradation in state-of-the-art tool agents when encountering non-collaborative users, as well as agent weaknesses under each non-collaborative condition such as escalated hallucinations and dialogue breakdowns. Our findings point to the need for methods that can improve agent robustness to the wide range of user behaviors encountered in deployment. We release the extensible simulation framework to help the community develop and stress-test tool agents under realistic conditions within their own service domains. Our code is available at https://github.com/holi-lab/NCUser.
@@ 1 Introduction
Tool agents engage in multi-turn dialogues where they interpret user requests, execute appropriate API calls, and communicate results back to users to complete specific tasks. Recent studies actively adopt user simulators to conduct multi-turn dialogue simulations between users and tool agents (Yao et al., 2024; Barres et al., 2025; Prabhakar et al., 2025). Unlike static dialogue datasets, these user simulators enable dynamic multi-turn interactions where the conversation flow adapts based on the agent’s responses and actions. This allows researchers to train and evaluate tool agents across diverse scenarios, capturing the interactive nature of real-world tool use. However, most existing user simulators and training datasets are agent-friendly, exhibiting only cooperative behaviors that fail to capture the complexity of real-world interactions, such as dealing with impatient users (Reynolds and Harris, 2009) or handling requests that are beyond the agent’s capabilities (Bitner et al., 1990). This hinders both developing robust agents and assessing their resilience to non-collaborative user behaviors in the real world.
@@
To address this limitation, we develop a novel user simulation framework that incorporates diverse non-collaborative behaviors observed in real-world interactions (Figure 1). We achieve this through two steps. First, we identify four types of non-collaborative user behaviors informed by marketing research, open-domain dialogue studies, and real-world user-agent interaction data: (1) Unavailable Services: users request functionalities beyond the agent’s API capabilities; (2) Tangential: users engage in free conversation unrelated to their primary task; (3) Impatience: users express frustration through emotional escalation when experiencing delays or service failures; and (4) Incomplete Utterances: users provide poorly articulated messages (§3.1). Second, we build a user simulator architecture that models these behaviors while ensuring goal-aligned simulation, i.e., reliably delivering all intents and information necessary to accomplish the given task (§3.2).
@@
In our main experiments, we leverage MultiWOZ (Budzianowski et al., 2018) and $\tau$-bench (Yao et al., 2024), to create stateful task-oriented dialogue environments and conduct comprehensive experiments to reveal the vulnerabilities of agents. We demonstrate that non-collaborative user behaviors lead to significant performance degradation for state-of-the-art LLMs. Our detailed analysis reveals how each behavior category impairs LLMs: unavailable service and incomplete utterance modes lead to immature tool utilization of the agents, while tangential and impatience modes expose shortcomings in their dialogue management. Moreover, when small LLMs are trained exclusively on typical (collaborative) scenarios, as many practitioners would do for the deployment of their services, the performance improvements on non-collaborative behaviors significantly lag behind those on collaborative behaviors. Lastly, we extend our non-collaborative user simulators to ColBench (Zhou et al., 2025) (which does not involve tool use) and MINT (Wang et al., 2024) (which involves user-agent collaboration) and observe disparate performance patterns across benchmarks. These results demonstrate that our extensible framework can preemptively diagnose potential agent weaknesses and yield insights across diverse domains. We hope that researchers working on LLM agents adapt our simulator to their own tasks for training and testing their agents against the complexities and challenges of real-world deployments.
@@
The main contributions of our work are summarized as follows: (1) We define four types of non-collaborative user behaviors and develop a user simulator framework that exhibits these behaviors while maintaining goal alignment. It outperforms simpler prompt-based approaches in creating challenging yet realistic dialogue scenarios. (2) We reveal that state-of-the-art tool agents exhibit significant performance degradation under non-collaborative conditions, providing insights into behavior-specific failure mechanisms that can guide the development of more robust agents. (3) We implement our user simulator on various settings, demonstrating the extensibility of our architecture and enabling researchers to evaluate and improve the agent resilience in real-world scenarios.
@@ 2 Related Works
It is important for tool agents to solve user requests in realistic scenarios. To evaluate these capabilities, several works have developed benchmarks (Qin et al., 2023; Patil et al., 2023; Tang et al., 2023) and simulation environments (Trivedi et al., 2024; Patil et al., 2025) to reflect realistic use cases. Despite their contributions, most of these works are limited to single-turn interactions, overlooking the fact that real-world task completion requires multi-turn conversations between users and agents. To address this gap, recent works such as Li et al. (2023) and Farn and Shin (2023) have generated multi-turn dialogues between users and tool agents, incorporating scenarios where agents engage in multi-turn conversations to elicit information from users (e.g., user intent, input parameter values). More complex scenarios include dialogues from Shim et al. (2025) where users fail to provide required parameters, and simulations from Laban et al. (2025) that model underspecification behavior. However, these works have still represented only interactions with collaborative users, with limited exploration of complicated situations arising from non-collaborative users. In our work, we develop a user simulator to reproduce these non-collaborative behaviors, thereby covering a broader spectrum of user-agent interactions.
@@
Adopting user simulators in dialogue simulation has been a well-established approach for developing task-oriented dialogue systems (Eckert et al., 1997; Sekulic et al., 2024; Schatztnann et al., 2005; Schatzmann et al., 2007), and has recently been adopted in tool agent research. The $\tau$-Bench series (Yao et al., 2024; Barres et al., 2025) and Apigen-mt (Prabhakar et al., 2025) propose prompt-based user simulators that incorporate the user goal, dialogue history, and instruction within the prompt to simulate goal-oriented conversations. To ensure goal-aligned behavior of user simulators, recent works have explored various techniques: Luo et al. (2024) employs LLM-based verifiers to inspect user utterances during dialogue simulation, while Mehri et al. (2025) trains user simulators through SFT and subsequent GRPO training with LLM-as-a-judge reward. However, existing user simulation research in tool agents has exclusively focused on collaborative user behavior, with no attempts to address or simulate the non-collaborative behaviors exhibited by real-world users. Furthermore, frameworks for building and evaluating goal-aligned user simulators remain heavily reliant on LLM-as-a-judge. To mitigate this, we investigate non-collaborative user behaviors through empirical studies and propose a novel user simulation method to reproduce these behaviors. We also enforce goal alignment, i.e., all goal-relevant information is delivered during the dialogue, and build a user simulator that satisfies it.
@@ 3.1 Non-Collaborative User Behaviors
We define our non-collaborative user behaviors grounded in marketing research and real-world user-LLM conversation studies. We first examined various types of user behaviors from both fields and clustered those that share similar features and then identified four representative categories: (1) Unavailable Services, (2) Tangential, (3) Impatience, and (4) Incomplete Utterances. We provide a detailed taxonomy and the underlying rationale for our behavior definitions in §A.
@@
This behavior stems from the user’s uncertainty about the agent’s boundaries, prompting them to request services beyond its capabilities (Bitner et al., 1990; Reynolds and Harris, 2005). We define this as requests that cannot be fulfilled using the available APIs—either because relevant APIs do not exist, or existing APIs lack the necessary parameters to accommodate the user’s specific requirements. For example, when a user requests “Book me a window seat on the train”, but the train booking API does not support seat selection options, the agent cannot fulfill the specific seat preference—making this an unavailable service request.
@@
Certain customers expect social rapport (Price et al., 1995; Beatty et al., 1996) or demand continuous attention (Jeung et al., 2018), and talkative users are known to be particularly challenging as they may become upset if not given space to express themselves (Geraghty, 2023; Jones, 2023; Coppell, 2023). We define tangential behavior as user utterances unrelated to goal completion—typically personal interests or background topics expressed through four dialogue acts adopted from open-domain dialogue research (Yu and Yu, 2021): (1) Factual Question, (2) Opinion Question, (3) General Opinion, and (4) Non-opinion Statement (See details in §C.2). We also model these users as raising complaints when agents ignore their tangential utterances. For example, a user booking a train might also ask, “By the way, where do you think I should visit first traveling in NA?”—a question unrelated to the booking task. When agents ignore such conversational attempts and focus solely on the transaction, users often feel unheard and express dissatisfaction.
@@
Users often exhibit impatience when experiencing service delays or failure notifications (Bitner et al., 1990; Xiao et al., 2022). We define impatience as emotional reactions triggered by agents taking excessive time to complete tasks or repeatedly failing to fulfill requests. Drawing from Reynolds and Harris (2009) and Harris and Reynolds (2004), we identify three dialogue acts that impatient users employ: (1) Belligerent Abuse, (2) Threat, and (3) Urge (See the details in §C.3). This manifests when, for example, an agent’s booking attempt fails repeatedly, prompting the user to shift from polite requests to aggressive demands: “Stop wasting my time and just get it done!”
@@
Following the principle of least effort (Zipf, 2016), users frequently produce incomplete or underspecified utterances (Laban et al., 2025) when expressing their intentions. We define incomplete utterances as poorly articulated messages. Real-world user-agent dialogue data from LMSYS (Zheng et al., 2024) and WildChat (Zhao et al., 2024) reveals two common patterns: extremely brief utterances, and prematurely sent messages. For example, when the user’s intended goal is “train reservation for 2 people”, a complete utterance is “I want to reserve a train for 2 people”. In contrast, incomplete utterances manifest as “Book train, 2” (extremely brief), or “I want to res” (prematurely sent).
@@ 3.2 Non-Collaborative User Simulation Framework
We construct a user simulator that addresses two key requirements: exhibiting non-collaborative behaviors while maintaining goal-aligned behavior (i.e., conveying intent and information from the user goal). Building on the collaborative user simulator from Yao et al. (2024) as our backbone, we enhance it with additional LLM-based modules to ensure goal-aligned behavior. Using this collaborative foundation, we incorporate the non-collaborative behaviors identified in §3.1 through various interventions with LLM modules (all prompts for LLM modules are in §G).
@@
As a foundation for non-collaborative simulation, we first establish a collaborative user simulator that communicates user intents and information to agents for goal completion. As illustrated in Figure 2, we adopt the simulation framework from Yao et al. (2024), which employs an LLM to generate contextually appropriate utterances based on three inputs: User Goal, Instruction, and Dialogue History, continuing until it generates an <END> token that terminates the dialogue. We implement this framework using GPT-4.1-mini as our user simulator.
To ensure goal-aligned behavior—where all intents and information specified in the user goal are communicated during dialogue—we enhance this base framework with two LLM-based modules inspired by Luo et al. (2024) (see User Simulator in Figure 2). First, following Laban et al. (2025), we shard user goals into information pieces and employ a dialogue state tracker to monitor which pieces have been conveyed at each turn. When the simulator attempts to terminate the dialogue without conveying all required information, the tracker ensures all remaining pieces are delivered. Second, we employ an ending verifier to prevent inappropriate dialogue termination (generating <END> token) even after all information has been delivered, particularly when the agent needs to execute actions or seeks user confirmation before proceeding. We provide all details including the base prompt, dialogue state tracker, information sharding, and ending verifier in §B.
@@
We simulate scenarios where users request services beyond the agent’s capabilities. As shown in Figure 3-a, we use GPT-4.1-mini to analyze the original user goal and identify potential services that would be unavailable to the agent. Based on this analysis, we generate three additional goal sentences that naturally extend the original goal while requiring either missing APIs or unsupported parameters. By concatenating with the original goal, we create augmented goals that enable the simulator to introduce unfulfillable requests during the dialogue (Examples in §C.1).
@@
We simulate users who introduce personal interests unrelated to their goals, expecting agents to acknowledge these tangential topics. Following the process in Figure 3-b, we generate tangential behavior through a two-stage approach. First, to ensure realistic and diverse tangential content, we randomly sample personas from Persona Hub (Ge et al., 2025). Using these persona characteristics, GPT-4o-mini generates tangential utterances that perform one of the four dialogue acts from §3.1. These tangential utterances are then merged with collaborative utterances from our base simulator to form the user’s complete utterance. Additionally, we simulate user dissatisfaction when these tangential utterances are ignored: GPT-4.1-mini checks whether the agent’s response addresses the tangential content, and if ignored, generates a complaint expressing disappointment that either replaces or augments the next collaborative utterance.
@@
We simulate impatient users who express anger when encountering agent failures or experiencing delays. Following the process in Figure 3-c, we trigger impatience utterance generator in two scenarios: (1) when the agent explicitly communicates failure or inability to complete a request, identified by GPT-4.1-mini, and (2) when the agent has not yet resolved the user’s goal despite the user having provided all required information (as determined by the dialogue state tracker), which we interpret as perceived delays. When either scenario is triggered, the system generates impatient utterances by randomly sampling from the three dialogue acts defined in §3.1. To model realistic anger escalation (Bushman, 2002), we implement a probabilistic activation mechanism where the likelihood of expressing anger increases with each triggering event, reflecting how real users become increasingly frustrated over time. Once anger is expressed, the user maintains a cynical utterance for all subsequent turns, reflecting persistent frustration after the initial outburst.
@@
We simulate users who produce incompletely written messages that complicate intent identification. Following the process in Figure 3-d, we model two types of utterance identified in §3.1: extremely brief utterances and prematurely sent messages. For extremely brief utterances, we perform style transfer on collaborative utterances using patterns collected from real user conversations in LMSYS (Zheng et al., 2024) and WildChat (Zhao et al., 2024), employing five randomly sampled examples for few-shot prompting (see the examples in §C.4). For prematurely sent messages, we simulate accidental mid-input truncation by randomly cutting collaborative utterances at various points. To ensure goal completion despite these truncations, the dialogue state tracker marks the truncated information as unsent, guaranteeing proper re-communication.